OpenStudy (anonymous):

Which of the following is a polynomial with roots -4, -2i, and 2i ?

OpenStudy (tamtoan):

if a polynomial with root a and b...then you can write it as (x-a)(x-b) = 0 , base on this, can you write your polynomial ?

OpenStudy (anonymous):

i don't really understand

OpenStudy (tamtoan):

example , x - 3 = 0, what is the root here ? ..i.e what is the value of x ?

OpenStudy (anonymous):

3

OpenStudy (tamtoan):

ok, so for example ...a polynomial of root 3 and 8 ...can you write this polonomial ?

OpenStudy (anonymous):

x-3=0; x-8=0?

OpenStudy (tamtoan):

that is not a polynomail ...should write it as (x-3).(x-8) = 0 ...this mean either x = 3 or x = 8 ..will make this polynomal equal to zero , now base on this, write your polynomial

OpenStudy (anonymous):

(x+4)(x+2i)(x-2i)=0?

OpenStudy (tamtoan):

ok, now do you know what i is? i.e ...definition of i

OpenStudy (anonymous):

equals -1 or something like that

OpenStudy (tamtoan):

no..i = sqrt(-1) , ie...i^2 = -1 ...also use this formula ...(x+a).(x-a) = x^2 - a^2 ...can you simply your polynomial there?

OpenStudy (anonymous):

no this confusing.

OpenStudy (tamtoan):

use this formula (x+a)(x-a) = x^2 - a^2 , tell me what is (x+2i).(x-2i) = ?

OpenStudy (anonymous):

required poly.. is (x+4)(x^2+4)=X^3+4x^2+4x +16

OpenStudy (tamtoan):

@matricked, giving the answer is not the way of helping him understand at all

OpenStudy (anonymous):

@tamtoan yup u r correct but if the person is cofusing then c

OpenStudy (tamtoan):

:) see his new question ? ,, it's basically the samething..give answer, he won't be able to do other exact same problem

OpenStudy (anonymous):

well i don't caree, and actually im a girl thank you :)